OpenStudy (anonymous):

i will fan and give medal If x+3/3 = y+2/2, then x/3= ?

hero (hero):

\[\frac{x + 3}{3} = \frac{y + 2}{2}\] right?

OpenStudy (anonymous):

yes

hero (hero):

Next, Cross multiply

OpenStudy (anonymous):

so i would do 2(x+3) and 3(y+2)?

hero (hero):

Next distribute

OpenStudy (anonymous):

2x+6 = 3y+6 next would i add the two 6 together and get 12?

hero (hero):

You would end up with 2x = 3y

hero (hero):

From there, you can manipulate the equation to figure out what x/3 equals

OpenStudy (anonymous):

y/2?

hero (hero):

Correct